"a movie theater charges 5.00 for an adult ticket and 2.00 for a childs ticket. One saturday the theater sold 785 ticket for 3280. How many childs tickets were sold for the movie that saturday?"
Answer by rfer(12678)

a+c=785 a=785-c
5a+2c=3280
--------------
5(785-c)+2c=3280
3925-5c+2c=3280
-3c=-645
c=215 children
a=570 adults